# Revisiting Differentially Private Hyper-parameter Tuning

### v2

@@ 4.6 Discussion
NTNV @@@@ $X=$ NTCV @@@@ $X=$ ETCV @@@@ $X=$
[…]
@@
TABLE III: Results summary for the lower bound $\varepsilon_{L}$ (averaged) obtained due to two different proxies under $\mathbf{z}=\mathbf{z}_{D}$ setup.
How does the score function (selection) leak privacy? This question is of much more interest than the previous two because it has never been studied in all previous auditing tasks (auditing the DP-SGD protocol) [39, 23, 37]. We summarise some representative results of our audit experiments in Table III. First, we observe that $\lambda_{a}$ and $\lambda_{b}$ have similar performance across all setups. Second, by comparing NTNV and NTCV, we see that manipulating the score function does not bring noticeable distinguishing advantages to the adversary.
This phenomenon may seem counter-intuitive and contrasts to the “magic” the adversary may expect: the adversary expects the (distribution of) score of the output when $\mathbf{z}$ is included in the training to be much different from that when $\mathbf{z}$ is not included; hence, manipulating the metric to select the best candidate, should make it much easier for the adversary to succeed the distinguishing game. However, our experimental results diverge from this expectation. The explanation lies in the following observations. $\lambda_{a}$ is computed based on the output resulting from selection by the score function, i.e., the selection action has already included the score information, indicating $\lambda_{b}$ is a duplicate of $\lambda_{a}$.
In principle, the score function’s output is differentially private as it is just post-processing of the trained model. With or without $\mathbf{z}$, the induced score distribution is close to each other, just as DP guarantees. Additional experimental results are presented in Appendix A.4 to validate such a claim. The “magic” cannot happen, and it holds for any score function, given that it is independent of the sensitive data.
5 Improved Privacy Results
In our previous empirical study, a conspicuous gap still exists between $\varepsilon_{U}$ derived by [42] and $\varepsilon_{L}$ obtained even under a strong adversary. Our study in the remaining sections shows such a gap exists for two reasons.
@@ 5 Improved Privacy Results
    Current generic privacy upper bound is not tight;
@@
    Adversary’s power is probably not strong enough because it is hard for the adversary to instantiate the worst-case score function $g$.
Regarding 1), we provide improved privacy results and elucidate on the special property of DP-SGD leading to the improvement; for 2), we present meaningful findings about the score function.
@@
[…]
@@ 5.1 Preliminaries: $f$-DP
The trade-off function governs the best one can achieve when distinguishing $P$ from $P^{\prime}$, i.e., by the optimal/smallest type II error ($\mathrm{FN}$) at fixed type I error ($\mathrm{FP}$). The optimal $\mathrm{FN}$ is achieved via the likelihood ratio test, which is also known as the fundamental Neyman–Pearson lemma [41] (please refer to Appendix A.1). We denote
[…]
@@
Algorithm 3 $f$-DP to $(\varepsilon,\delta)$-DP
1:trade-off function $f$, $\delta$ with $\delta\geq 1-f(0)$
2:Compute $\varepsilon=\inf\{a:f(x)\geq 1-\delta-\mathrm{e}^{a}x,\forall x\in[0,1]\}$ via binary search
3:$\max\{0,\varepsilon\}$
@@
The analytical expression of $\mu$-GDP is determined by applying the Neyman–Pearson lemma on the hypothesis testing problem of distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$ [17]. $\mathcal{M}$ satisfying $\mu$-GDP means that distinguishing $\mathcal{M}(X)$ from $\mathcal{M}(X^{\prime})$ is at least as hard as distinguishing two univariate Gaussians $\mathcal{N}(0,1),\mathcal{N}(\mu,1)$ based on a single draw. Figure 6 explains why $(\varepsilon,\delta)$-DP is loose: $(\varepsilon,\delta)$-DP is strictly more conservative than $\mu$-GDP when characterizing the privacy of Gaussian mechanism.
@@
DP-SGD is asymptotically $\mu$-GDP. The pivotal role of $\mu$-GDP is that it asymptotically characterizes the privacy of any DP-SGD instance having many compositions of Gaussian mechanisms [17], as shown in the following.
@@
5.2 Our Contribution: Improved Results
For Corollary 2, the error (pointwise error between the asymptotical GDP trade-off function and the true trade-off function) decays at a rate of $1/\sqrt{N}$ for DP-SGD, as shown by the analysis in [17]. Therefore, using Corollary 2 requires $N$ to be large enough. Such a condition holds for probably most DP-SGD applications, especially for training large models (e.g., $N>10^{4}$ in [2] and $N>10^{5}$ in [27, 51]). Based on all of the above preparations, we are ready to approach our privacy problem by filling in the missing part of Equation (11) based on Corollary 2:
@@ 5.2 Our Contribution: Improved Results
[…]
@@
Recall in Section 4.6, we showed that the adversary gains no advantage even if the score function $g$ is maliciously manipulated. A natural question arises: is there any score function that brings more advantage to the adversary? The answer is positive as there is a special type of score function that tends to leak more sensitive information, regardless of whether manipulated or not.
One-to-one mapping $g$ is the worst-case necessarily. Before introducing our main privacy result, we have a final question remaining to be answered: how does the score function $g$ affect the privacy of $\mathcal{H}$? Specifically, as a function that maps the output of the base algorithm to a real number, If some $g$ happens to map two distinct inputs to the same score (hence, a randomized tie-breaking will be enforced), how does such $g$ affect the privacy of $\mathcal{H}$ compared to one-to-one mapping score functions?
@@
Theorem 2 (Necessary worst-case $g$, proof in Appendix B.2).
Let distribution $P$ be over some finite alphabets $\Gamma$, and define a distribution $F_{k,g}$ as follows. First, make $k>0$ independent samples $\{x_{1},x_{2},\cdots,x_{k}\}$ from $P$; second, output $x_{i}$ such that the score $g(x_{i})$ computed by a score function $g:\Gamma\rightarrow\mathbb{R}$ is the maximum over these samples. Similarly, we define another distribution $P^{\prime}$ over the same alphabets $\Gamma$ and derive a distribution $F^{\prime}_{k,q}$ as the counterpart to $F_{k,g}$.
@@
The above result is derived under RDP (Definition 2) due to analytical feasibility, but it tells us crucial facts: A score function that induces a strict total order for elements in $\Gamma$ tends to be less private. Thus, a one-to-one mapping is necessary to be the worst case for the score function $g$. Theorem 2 also holds when $\Gamma$ is infinite because Rényi divergence can be approximated arbitrarily well by finite partition [50, Theorem 10].
With $g$’s necessary condition determined, we formalize some notation and introduce our improved privacy results in the following.
@@
Theorem 3 (General form, proof in Appendix B.3).
@@
Algorithm 4 $f$-DP Accountant for $\mathcal{H}$
@@
2:$\delta\leftarrow\delta_{\mathcal{H}}/\omega_{\xi}(1)$$\triangleright$ $\omega_{\xi}$ is from Equation (16)
3:$\varepsilon\leftarrow$ input $f$ and $\delta$ to Algorithm 3
4:$\varepsilon_{\mathcal{H}}\leftarrow\varepsilon+\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f(a))}}}$
5:$\varepsilon_{\mathcal{H}}$
@@
Why we can obtain improvements. In the following, we show why modeling the base algorithm with $f$-DP other than $(\varepsilon,\delta)$-DP brings improvement. By post-processing property [17], the score outputted by the score function satisfies the same $f$-DP as that of the base algorithm. Term $a$ in Equation (18) can be seen as the $\mathrm{FN}$ for distinguishing $P$ from $P^{\prime}$ and $b=f(a)$ is the optimal $\mathrm{FP}$ at $\mathrm{FN}=a$.
For DP-SGD instance satisfying $\mu$-GDP, we have $b=G_{\mu}(a)$, however, if its privacy is represented by some $(\varepsilon,\delta)$-DP (or $f_{\varepsilon,\delta}$-DP) we can only get $b=f_{\varepsilon,\delta}(a)$ in Equation (18). Note that figure 6 tells us that $G_{\mu}(a)\geq f_{\varepsilon,\delta}(a)$, hence modeling the base algorithm by $(\varepsilon,\delta)$-DP only leads to higher (weaker) $\varepsilon_{\mathcal{H}}$ value in Equation (18) as $\omega_{\xi}$ is increasing. The following numerical example demonstrates the superiority of modeling within the $f$-DP framework.
@@
For $\delta=10^{-5}$, the base algorithm is also $(4.36,10^{-5})$-DP or $f_{4.36,10^{-5}}$-DP. If $b=G_{1}(a)$ in Equation (18), which is how we represent the base algorithm’s privacy, we have $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(G_{1}(a))}}}=3.3$; however, if $b=f_{4.36,10^{-5}}(a)$, which equals to how the base algorithm is modeled by Liu et al. [30], we can only have $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f_{4.36,10^{-5}}(a))}}}=16.5>3.3$. Thus, a huge improvement is obtained, and this is due to the saved shaded area in gray shown in Figure 6.
5.3 Discussion
@@
Improvement is also possible in general. Our above example shows that representing the privacy of the base algorithm with finer resolution (from $(\varepsilon,\delta)$-DP to $f$-DP) leads to improvements in the privacy result for the private selection problem. Similar conclusions also hold when switching from RDP [42] to $f$-DP as RDP is also observed to be lossy within the $f$-DP framework [58], i.e., RDP shares the same weakness as that of the $(\varepsilon,\delta)$-DP. Therefore, for any other private mechanisms analyzed by $(\varepsilon,\delta)$-DP or RDP, switching to $f$-DP also brings improvements possibly.
Our results are both tighter and more generalizable. As will be shown in Section 6.2, our result is tighter. Moreover, consider a $\xi$ distribution such that $\operatorname{Pr}_{\xi}(1)=1$ (i.e., run the base algorithm only once), we see that Equation (18) becomes $\varepsilon_{\mathcal{H}}=\varepsilon$ and $\delta_{\mathcal{H}}=\delta$ for any $(\varepsilon,\delta)$ pair, which is the base algorithm’s privacy guarantee. This shows that our result is tight for general distribution $\xi$ in terms of how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies. In contrast, analysis by RDP can’t achieve such tightness as converting RDP to $(\varepsilon,\delta)$-DP is lossy [58].
As for the generalizability, we note that 1) [30] only works with $\xi$ being the geometric distribution; 2) [42] only works with $\xi$ being the TNB family and Poisson distribution. It is unknown how to handle the case where $\xi$ is some arbitrary distribution, and it seems like at least case-by-case analysis is needed if using previous approaches. In contrast, our result applies to any distribution $\xi$ of protocol $\mathcal{H}$. Also, note that computing $\omega_{\xi}$ for any reasonable distribution is always numerically stable as $\omega_{\xi}(x)$ is bounded in $[0,1]$.
6 Further Evaluation
Unlike the experiments in Section 4, which relies on some ad hoc audit setups, our further audit experiments in this section make reductions to the audit and confine to the idealized settings of the threat model of DP. We aim to explore the tightness of our improved result by finding where the privacy lower bound is under an adversary with compelling and theoretically justified power.
6.1 Stronger Audit via Reduction
1) Base algorithm reduction. To have stronger privacy audit result over $\mathcal{H}$, it is imperative to first have stronger audit on $\mathcal{H}$’s the base DP-SGD algorithm. This part is devoted to this topic.
@@ 6.1 Stronger Audit via Reduction
Our threat model will be based on the assumption made by DP, i.e., the adversary knows all but one data, a.k.a. the strong adversary assumption [13, 29]. Translated to our setting, the adversary knows the membership of all data used to update the model in each iteration except for the differring data $\mathbf{z}$. This means the adversary can always subtract the gradient of other data from $p_{i}$ in each iteration. Based on such setups, any adjacent dataset pair $X,X^{\prime}$ is equivalent to $X=\emptyset,X^{\prime}=\mathbf{z}$ from the adversary’s view, i.e., we audit in an idealized setting.
The remaining part is how we form a compelling proxy and let the adversary make assertions based on it. This is done via two reductions for the base algorithm (DP-SGD).
@@
This is because a noise vector $R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ and its rotated version $U_{\mathbf{z}}R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ possessing $\sigma^{2}\mathbb{I}^{d}$ covariance matrix are all coordinate-wise independent. To serve the distinguishing purpose: $\mathbf{z}$ was/was not used, it suffices to characterize the private gradient $p_{i}$ by $\bar{p}_{i}$ as a univariate random variable (the first coordinate) for the distinguishing purpose as follows.
@@
Therefore, our above analysis justifies the sufficiency of the Dirac canary $\mathbf{z}=\mathbf{Z}_{D}$ being worst-case in our previous experiments in Section 4.5.
❷ Second reduction (from Equation (22) to Equation (24)). By the first reduction and from the adversary’s view, the DP-SGD’s output $\bar{y}=\{\bar{p}_{1},\bar{p}_{2},\cdots,\bar{p}_{N}\}$ is essentially an observation of $N$ i.i.d. samples from $\mathcal{N}(a,C^{2}\sigma^{2})$ where $a$ is either 00 or $C$. Recall the adversary’s goal is to distinguish $X$ or $X^{\prime}$ was used; this is equivalent to determining $a=0$ or $a=C$.
As both distributions in Equation (22) are Gaussian, we can use the sufficient statistics for estimating $a$ [21, 12], which is the mean of such $N$ i.i.d. samples: $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}\bar{p}_{i}$. Sufficient statistics do not lose any information for estimating $a$.
Finally, we can reduce the privacy of the base algorithm to an equivalent game for the adversary as
@@
Figure 7: The distinguishing game for $\mathcal{H}$ with the base algorithm reduced. Two univariate Gaussians $P=\mathcal{N}(0,1)$ and $P^{\prime}=\mathcal{N}(\frac{\sum_{i=1}^{N}b_{i}}{N}\sqrt{N}/\sigma,1)$ correspond to adjacent dataset pair $X,X^{\prime}$. At each execution, the trainer chooses one distribution and feeds it to $\mathcal{H}$; then an integer $k>0$ is sampled from $\xi$ followed by independently sampling $k$ times from $\hat{P}$. After getting such $k$ samples, the best sample $\bar{y}_{i}$ is selected based on a score function $g(x)=x$. The adversary then makes binary assertions by 1) simply taking the returned best sample as the proxy and 2) comparing it to some threshold. The game $\mathcal{G}$ is simulated $10^{7}$ times.

### v3

@@ IV-F Phenomenon Highlight
NTNV @@ $X=$ NTCV @@ $X=$ ETCV @@ $X=$
$\mathbb{F}$ $\mathbb{C}$ $\mathbb{S}$ $\mathbb{F}$ $\mathbb{C}$ $\mathbb{S}$ $\emptyset$
@@
TABLE III: Results summary for privacy audit result $\varepsilon_{L}$ (averaged over 5 runs) obtained under two different proxies when the canary $\mathbf{z}=\mathbf{z}_{D}$. We also highlight the base algorithm’s upper bound $\varepsilon_{B}$ and upper bound for hyper-parameter tuning using [42]’s method.
How does the score function (selection) leak privacy? This is a key question. Unlike prior auditing works [39, 23, 37], it has not been studied. Table III summarizes key audit results. First, $\lambda_{a}$ and $\lambda_{b}$ perform similarly across all settings. Second, comparing NTNV and NTCV shows that manipulating the score function offers no clear advantage to the adversary.
This may seem surprising. One might expect that tailoring the metric to pick the best candidate would make distinguishing easier. But our results show otherwise. This is because $\lambda_{a}$ already reflects the effect of the score function—it is computed from the selected output. So $\lambda_{b}$ adds no new information.
In fact, the score function output is post-processing of a DP algorithm, and thus remains differentially private. Whether or not $\mathbf{z}$ is included, the induced score distributions are close—just as DP guarantees. So no “magic” advantage exists, regardless of the score function, if it’s independent of sensitive data.
Is it safe to tune hyper-parameter while only assuming a single run’s privacy cost? Our weakest audit setup, which mimics practical use, suggests yes. Tuning adds little to no extra leakage beyond the base algorithm. This supports the common—but non-rigorous—practice of ignoring tuning overhead before private selection methods were formalized [30, 42].
However, under our strongest audit, tuning does leak more privacy. This confirms the importance of formally accounting for the tuning step, justifying both the previous work and our improved bound in Section V.
Can we get a higher lower bound on any other datasets than what we get in ETCV setting? The ETCV setup already achieves dataset-independent lower bounds. Since the adversary can inject arbitrary gradient canaries, any bound achievable on real data is also achievable in this setting.
V Improved Theoretical Results
In the previous empirical study, a conspicuous gap exists between $\varepsilon_{U}$ derived by [42] and $\varepsilon_{L}$, this makes it interesting to investigate the reason behind it. Our study in the remaining sections shows that such a gap exists for two reasons.
@@ V Improved Theoretical Results
    Current generic privacy upper bound is not tight for DP-SGD;
@@
    Adversary’s power is not strong enough because it is hard for the adversary to instantiate the worst-case score function $g$.
Regarding 1), we provide improved privacy results and elucidate on the special property of DP-SGD leading to the improvement; Our analysis is generalizable beyond DP-SGD, i.e., as will be shown, our analysis works for any base algorithm that can be expressed within the $f$-DP framework. For 2), we present meaningful findings about the score function.
@@
subject to:   $\text{\quad$\mathcal{H}$ satisfies $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP given $\delta_{\mathcal{H}}$};$
@@ V-A Preliminaries: $f$-DP
The trade-off function governs the best one can achieve when distinguishing $P$ from $P^{\prime}$, i.e., by the optimal/smallest type II error ($\mathrm{FN}$) at fixed type I error ($\mathrm{FP}$). The optimal $\mathrm{FN}$ is achieved via the likelihood ratio test, which is also known as the fundamental Neyman–Pearson lemma [41] (please refer to Appendix A-A). We denote
$\text{$g\geq f$ if $g(x)\geq f(x),\forall x\in[0,1]$}.$
@@
The analytical expression of $\mu$-GDP is due to the application of Neyman–Pearson lemma on distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$ [17]. $\mathcal{M}$ satisfying $\mu$-GDP means that distinguishing $\mathcal{M}(X)$ from $\mathcal{M}(X^{\prime})$ is at least as hard as distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$. Figure 5 explains why $(\varepsilon,\delta)$-DP is loose: $(\varepsilon,\delta)$-DP is strictly more conservative than $\mu$-GDP when characterizing the privacy of Gaussian mechanism.
@@
Algorithm 3 ff-DP to $(\varepsilon,\delta)$-DP [17]
1:$f$, trade-off function; $\delta$, privacy parameter
2:If $\delta<1-f(0)$, return $\infty$
3:Compute $\varepsilon=\inf\{a:f(x)\geq 1-\delta-\mathrm{e}^{a}x,\forall x\in[0,1]\}$ via binary search
[…]
@@
DP-SGD is asymptotic $\mu$-GDP. $\mu$-GDP is pivotal because it asymptotically characterizes the privacy of any DP-SGD instance having many compositions of Gaussian mechanisms [17]:
@@
V-B Our Contribution: Improved Results
A critical fact about Corollary 2 is that computing the exact trade-off function is #P-hard [17] (even harder than NP problems), which makes it necessary to resort to approximations if one aims at tighter results within the $f$-DP framework. Specifically, the error (pointwise error between the asymptotical GDP trade-off function and the true trade-off function) decays at a rate of $1/\sqrt{N}$ for DP-SGD, shown by [17]. Therefore, using Corollary 2 requires $N$ to be large enough. Such a condition holds for probably most DP-SGD applications, especially for training large models (e.g., $N>10^{4}$ in [2] and $N>10^{5}$ in [27, 51]).
Based on all of the above preparations, we are ready to approach our privacy problem by filling in the missing part of Equation (12) based on Corollary 2:
@@ V-B Our Contribution: Improved Results
subject to:   $\text{\quad$\mathcal{H}$ satisfies $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP given $\delta_{\mathcal{H}}$};$
@@
In Section IV-F, we showed that the adversary gains no advantage even if the score function $g$ is maliciously manipulated. A natural question arises: is there any score function that brings more advantage to the adversary?
One-to-one mapping $g$ is the worst-case necessarily. The score function $g$ is a function that maps the output of the base algorithm to a real number, If some $g$ happens to map two distinct inputs to the same score (hence, a randomized tie-breaking will be enforced), how does such $g$ affect the privacy of $\mathcal{H}$ compared to one-to-one mapping score functions?
@@
Theorem 2 (Necessary worst-case $g$, proof in Appendix B-B).
Let distribution $P$ be over some finite alphabets $\Gamma$, and define a distribution $F_{k,g}$ as follows.
First, make $k>0$ independent samples $\{x_{1},x_{2},\cdots,x_{k}\}$ from $P$; second, output $x_{i}$ such that the score $g(x_{i})$ computed by a score function $g:\Gamma\rightarrow\mathbb{R}$ is the maximum over these samples. Similarly, we define another distribution $P^{\prime}$ over the same alphabets $\Gamma$ and derive a distribution $F^{\prime}_{k,q}$ as the counterpart to $F_{k,g}$.
@@
The above result is derived under RDP (Definition 2) and it tells us crucial facts: A score function that induces a strict total order for elements in $\Gamma$ tends to be less private. Thus, a one-to-one mapping is necessary to be the worst case for the score function $g$.
Note that, in previous work [30, 42], the score function $g$ is assumed to be one-to-one mapping by default for simplicity. We show that such treatment is valid due to privacy considerations; to our knowledge, the above theorem is the first rigorous proof validating such an assumption.
Theorem 2 also holds when $\Gamma$ is infinite because Rényi divergence can be approximated arbitrarily well by finite partition [50, Theorem 10]. With $g$’s necessary condition determined, we can introduce our improved privacy results.
@@
Theorem 3 (General form, proof in Appendix B-C).
@@
Algorithm 4 ff-DP Accountant for $\mathcal{H}$
@@
[…]
@@
Intuitive explanation of our results. The intuition is that the selection (choosing the best of many independent runs) results in a different output distribution than when running the base algorithm only once. And it will deteriorate the final privacy bound, this is shown in Equation 19: there is an increase (deteriorating) of the $\varepsilon$ parameter compared to the base algorithm’s parameter. How the results deteriorate depends on $\xi$.
Modeling the base algorithm with $f$-DP instead of $(\varepsilon,\delta)$-DP leads to tighter bounds. By the post-processing property [17], the score output retains the same $f$-DP as the base algorithm. In Equation (19), $a$ represents the false negative (FN), and $b=f(a)$ is the optimal false positive (FP) at that FN.
If the base algorithm satisfies $\mu$-GDP, then $b=G_{\mu}(a)$. But if modeled using $(\varepsilon,\delta)$-DP, we only get $b=f_{\varepsilon,\delta}(a)$. Figure 5 shows that $G_{\mu}(a)\geq f_{\varepsilon,\delta}(a)$. Since $\omega_{\xi}$ is increasing, this gives a tighter (smaller) $\varepsilon_{\mathcal{H}}$ when using GDP. The following example illustrates the gain from using $f$-DP.
@@
For $\delta=10^{-5}$, the base algorithm is also $(4.36,10^{-5})$-DP or $f_{4.36,10^{-5}}$-DP. If $b=G_{1}(a)$ in Equation (19), which is how we represent the base algorithm’s privacy, we have $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(G_{1}(a))}}}=3.3$; however, if $b=f_{4.36,10^{-5}}(a)$, which equals to how the base algorithm is modeled by Liu et al. [30], $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f_{4.36,10^{-5}}(a))}}}=16.5>3.3$ is only what we can derive. Thus, a huge improvement is obtained, and this is due to the saved shaded area in gray shown in Figure 5.
V-C Significance Statement
Our result is generalizable and tighter. We observe that 1) [30] only supports geometric $\xi$, and 2) [42] only supports truncated negative binomial and Poisson $\xi$. It is unclear how to handle arbitrary $\xi$, and prior results require manual, case-specific analysis. In contrast, our result works for any $\xi$ in protocol $\mathcal{H}$. Computing $\omega_{\xi}$ is always numerically stable, as $\omega_{\xi}(x)$ is bounded on $[0,1]$.
As shown in Section VI-B, our bound improves over prior work. For example, if $\xi$ always outputs 1 (i.e., run the base algorithm once), Equation (19) gives $\varepsilon_{\mathcal{H}}=\varepsilon$ and $\delta_{\mathcal{H}}=\delta$, matching the base guarantee. This shows our result is tight for general $\xi$. In contrast, RDP-based analysis is loose due to lossy conversion to $(\varepsilon,\delta)$-DP [58].
@@ V-C Significance Statement
Extension beyond DP-SGD. Our above example shows that representing the privacy of the base algorithm with finer resolution (from $(\varepsilon,\delta)$-DP to $f$-DP) leads to improvements in the privacy upper bound. Similar conclusions also hold when switching from RDP [42] to $f$-DP as RDP is also observed to be lossy within the $f$-DP framework [58], i.e., RDP shares the same weakness as that of the $(\varepsilon,\delta)$-DP. We select DP-SGD as our base algorithm because of its popularity in the literature, but our result is not limited to DP-SGD. In fact, any private base algorithm analyzed by $(\varepsilon,\delta)$-DP or RDP can be represented by $f$-DP with finer resolution. And switching to $f$-DP and using our privacy accountant can also bring improvements. The reason is depicted in Figure 5: using $f$-DP avoids the unnecessary region shaded in gray.
VI Further Evaluation
VI-A Stronger Audit via Reduction
Motivation for final audit trial. In the presence of our improved privacy upper bound, we immediately want to assess its tightness by privacy audit for general $\xi$ such that $\operatorname{Pr}_{\xi}(1)<1$. This requires we derive reasonably strong lower bounds to be informative. We should avoid ad hoc audit setups for real-world training tasks (Section IV). We need to form our audit with theoretical-justified power.
This section is to serve such a purpose. A part of the design considerations relies on our Theorem 2 in the last section.
@@ VI-A Stronger Audit via Reduction
1) Base algorithm reduction. Our threat model will be based on the assumption made by DP, i.e., the adversary knows the membership of all data used to update the model in each iteration except for the differring data $\mathbf{z}$ (the strong adversary assumption [12, 29]).
This means the adversary can always subtract the gradient of other data from $p_{i}$ in each iteration. Hence, any adjacent dataset pair $X,X^{\prime}$ is equivalent to $X=\emptyset,X^{\prime}=\mathbf{z}$ from the adversary’s view. This allows us to make two reduction steps for the base algorithm (DP-SGD).
@@
Because a noise vector $R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ and its rotated version $U_{\mathbf{z}}R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ possessing $\sigma^{2}\mathbb{I}^{d}$ covariance matrix are all coordinate-wise independent. To serve the distinguishing purpose: $\mathbf{z}$ was/was not used, it suffices to characterize the private gradient $p_{i}$ by $\bar{p}_{i}$ as a univariate random variable (the first coordinate) for the distinguishing purpose as follows.
@@
❷ Second reduction (from Equation (23) to Equation (25)). Base on previous reduction, the DP-SGD’s output $\bar{y}=\{\bar{p}_{1},\bar{p}_{2},\cdots,\bar{p}_{N}\}$ is essentially an observation of $N$ i.i.d. samples from $\mathcal{N}(a,C^{2}\sigma^{2})$ where $a$ is either 0 or $C$. Recall the adversary’s goal is to distinguish $X$ or $X^{\prime}$ was used; this is equivalent to determining $a=0$ or $a=C$.
As both distributions in Equation (23) are Gaussian, we can use the sufficient statistics for estimating $a$ [21, 11], which is the mean: $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}\bar{p}_{i}$. Sufficient statistics do not lose any information for estimating $a$. Finally, we can reduce the privacy of the base algorithm to an equivalent game for the adversary as
@@
2) Instantiate the score function. Based on our above reduction, to serve the distinguishing purpose, a model obtained by DP-SGD can be “treated” as a real number sampled from univariate Gaussian or its shifted counterpart corresponding to $X$ or $X^{\prime}$ was used. The order induced by the score function $g$ is now over $\mathbb{R}$. To have stronger audit results in our idealized attack, we need to instantiate the worst-case score function, and our Theorem 2 tells us one-to-one mapping score function $g$ is the worst case necessarily.
However, Theorem 2 remains silent on the specific analytical form of $g$ in the worst case. There can be infinitely many one-to-one mapping functions $\mathbb{R}\rightarrow\mathbb{R}$; for implementation purposes, we now fix a score function $g(x)=x$, i.e., we take $g$ is strictly increasing and note that all strictly increasing functions induces the same order over $\mathbb{R}$ regardless of its analytical form. Finally, we present the distinguishing game of our reduced case in Figure 6, which will be simulated many times, allowing high-confident conclusions of the lower bound following Section II-B.
Figure 6: The distinguishing game for $\mathcal{H}$ by reduction. $P,P^{\prime}$ are two univariate Gaussians shown in Equation (27), corresponding to adjacent dataset pair $X,X^{\prime}$. The adversary then makes binary assertions by comparing the best sample to some threshold. The game $\mathcal{G}$ is simulated $10^{7}$ times.
VI-B Privacy Results Comparison
Privacy results. We present the lower bound $\varepsilon_{L}$, our improved privacy results $\varepsilon^{O}_{\mathcal{H}}$ and previous privacy results $\varepsilon_{\mathcal{H}}^{P}$ [42] in Table IV. We can see that the audit is rather powerful: $\varepsilon_{L}$ $\varepsilon_{B}$ with a noticeable margin, once again confirming the very action of selecting the best does leak additional privacy beyond the base algorithm’s privacy budget. On the other hand, our privacy result $\varepsilon_{\mathcal{H}}^{O}$ is better than the previous result $\varepsilon_{\mathcal{H}}^{P}$ under all setups.
Implication. The improvement we obtain is much more significant than the $\epsilon^{O}_{\mathcal{H}}$ value shows cosmetically. For instance, the results due to our improved upper bound $\epsilon^{O}_{\mathcal{H}}$ under $(\eta=1,\nu=10^{-3})$ are even consistently smaller than $\varepsilon_{\mathcal{H}}^{P}$ of previous upper bound under $(\eta=1,\nu=10^{-2})$. Thus, our improved analysis can allow trying significantly more hyper-parameter candidates while even consuming less privacy budget. For private hyper-parameter tuning applications, this translates to improved utility of the trained model for free.
To show the downstream gain using our method, we present the privacy result comparison in Table V. As can be seen from Table V, under the same privacy constraint, we allow more running (in expectation) hyper-parameter candidates, and the final achievable testing accuracy is consistently better as expected.